### v2

# Identification of generic finite metric spaces by eigenvalues

Abstract
Similarity matrices are fundamentals of the theory of magnitude. We show that finite metric spaces can be identified by the characteristic polynomials of the similarity matrices of $tX$ $(t>0)$, where $tX$ is a homothetical expansion of $X$ by a factor $t$, if we restrict ourselves to a certain open dense subset of the set of finite metric spaces.
Keywords: Similarity matrix, magnitude
2020 Mathematics Subject Classification: 51F99
1 Introduction
[…]
Let $tX$ $(t>0)$ be a metric space obtained by a homothetical expansion of $X$ by a factor $t$, $tX=(X,td)$. Then $Z_{tX}$ is invertible for all but finitely many $t>0$ ([4] Proposition 2.2.6 i). The map $t\mapsto|tX|$ is called the magnitude function, which is a partially defined function on $(0,+\infty)$. Let us denote it by $M_{X}(t)$. By putting $q=e^{-t}$ we obtain the formal magnitude, which is a generalized formal power series allowing non-integer powers of $q$, denoted by $m_{X}(q)$.
The magnitude homology $MH^{l}_{\ast}$ was introduced for graphs by Hepworth and Willerton [3] and for enriched categories by Leinster and Shulman [6]. The definition by the former can easily be generalized to finite metric spaces. The formal magnitude can be considered as the Euler characteristic of the magnitude homology ([3] Theorem 2.8222It was proved for graphs in [3], the proof works for finite metric spaces.).
Let us consider a question of to what extent the magnitude can distinguish finite metric spaces. We showed in [8] that a generic finite metric space can be identified by the formal magnitude in the set of generic finite metric spaces, where we say that a finite metric space $X=\{x_{1},\dots,x_{n}\}$ is generic if $d(x_{i},x_{j})$ $(i<j)$ are rationally independent. The genericity condition was introduced to prevent data from getting mixed, resulting in the desired information being lost. On the other hand, there are known examples of spaces that are not isometric but have the same magnitude or even isomorphic magnitude homology groups (Section 3).
In this paper we weaken the condition of genericity to have the notion of weakly $3$-genericity which gives an open dense subset of the set of finite metric spaces in the sense of Roff-Yoshinaga’s parametrization of finite metric spaces ([10]), and show that in this subset spaces can be constructed from the characteristic polynomials or equivalently the eigenvalues of the similarity matrix of $tX$ $(t>0)$. In other words, the characteristic polynomial $P_{X}(t,\lambda)=\det(\lambda I-Z_{tX})$, considered as a function with two variables, is the complete invariant for weakly $3$-generic finite metric spaces.
One of the advantages of considering the characteristic polynomials is that they seem to be good at identifying spaces. In fact we show that four-point metric spaces can be identified by the characteristic polynomial, in contrast to the fact that there are four-point sets with the same magnitude that are not isometric. In Section 3 we listed some known examples of finite metric spaces that are not isometric but have the same magnitude or even isomorphic magnitude homology groups. In each case, the characteristic polynomials are different. Although, the characteristic polynomials of the similarity matrices are not the complete invariant of finite metric spaces as illustrated in Section 4.
The author used Maple 2023 for the computations in Subsection 2.2 and Section 3. To make the expressions easier to read, we use $q=e^{-t}$ as the variable. Throughout the article, a (metric) space means a finite metric space.
Acknowledgement: The author thanks Yu Tajima deeply for the information on Whitney twist.
2 Identification by eigenvalues of similarity matrix
Let $X=\{x_{1},\dots,x_{n}\}$ and put $d_{ij}=d(x_{i},x_{j})$, which we call edge length. A sequence $\Gamma=(x_{i_{0}},\dots,x_{i_{k}})$ of $k+1$ points of $X$ such that adjacent points are different is called a $k$-path of $X$. Its length is defined by the sum of $k$ edge lengths; $\ell(\Gamma)=\sum_{j=1}^{k}d_{i_{j-1}i_{j}}$.
Definition 2.1
Let $k\geq 2$. Bt a simple closed $k$-path or a single $k$-cycle we mean the equivalence class of sequences of $k$ mutually distinct points $\Gamma=[x_{i_{1}},\dots,x_{i_{k}}]=(x_{i_{1}},\dots,x_{i_{k}})/\sim$, where the equivalence is generated by the cyclic shift;
$(x_{i_{1}},x_{i_{2}},\dots,x_{i_{k}})\sim(x_{i_{2}},\dots,x_{i_{k}},x_{i_{1}}).$
The length of a simple closed $k$-path is defined to be the sum of $k$ edge lengths; $\ell(\Gamma)=\sum_{j=1}^{k-1}d_{i_{j}i_{j+1}}+d_{i_{k}i_{1}}.$
By a $k$-cycle we mean a disjoint union of single cycles such that the sum of the numbers of vertices is equal to $k$. Its length is defined to be the sum of the lengths of constituent single cycles. Let $c(\Gamma)$ be the number of connected components of a $k$-cycle $\Gamma$. Let $\mathcal{C}_{k}(X)$ be the set of $k$-cycles of $X$. Put $\mathcal{C}_{1}(X)=\emptyset$.
Remark that both $k$-paths and $k$-cycles are oriented.
By putting $q=e^{-t}$ as in [5], the similarity matrix is expressed by $z_{X}(q)=\left(q^{d_{ij}}\right)_{i,j}$. Let $p_{X}(q;\lambda)$ be the characteristic polynomial of $z_{X}(q)$; $p_{X}(q;\lambda)=\left|\lambda I-z_{X}(q)\right|$. Each coefficient of $\lambda^{k}$ is a “generalized polynomial ” of $q$ that allow non-integer exponents. Let $\lambda_{1}(q),\dots,\lambda_{n}(q)$ be the eigenvalues of $z_{X}(q)$. Let $\bar{p}_{X}(q;\mu)$ be a polynomial obtained from $p_{X}(q;\lambda)$ by setting $\mu=\lambda-1$; $\bar{p}_{X}(q;\mu)=p_{X}(q;\mu+1)$. Put $\mu_{i}(q)=\lambda_{i}(q)-1$, and let $\tau_{k}(q)$ $(1\leq k\leq n)$ be the $k$-th elementary symmetric polynomial of $\mu_{1}(q),\dots,\mu_{n}(q)$.
Lemma 2.2
There holds
$\tau_{k}(q)=\sum_{\Gamma\in\mathcal{C}_{k}(X)}{(-1)}^{c(\Gamma)+k}q^{\,\ell(\Gamma)}\qquad(1\leq k\leq n),$ (2.1)
where $c(\Gamma)$ is the number of connected components of $\Gamma$.
Proof. When $k=1$, $\tau_{1}(q)={\rm trace}\,(I-z_{X}(q))=0$.
Suppose $k\geq 2$. As there holds
$\bar{p}_{X}(q;\mu)=\prod_{i=1}^{n}(\mu-\mu_{i}(q))=\mu^{n}+\sum_{l=1}^{n}(-1)^{l}\tau_{l}(q)\,\mu^{n-l}=\left|\begin{array}[]{cccc}\mu&-q^{d_{12}}&\cdots&-q^{d_{1n}}\\ -q^{d_{21}}&\mu&&\vdots\\ \vdots&&\ddots&-q^{d_{n-1\,n}}\\[2.84526pt] -q^{d_{n1}}&\cdots&-q^{d_{n\,n-1}}&\mu\end{array}\right|,$ (2.2)
the coefficient of $\mu^{n-k}$ is given by
$\sum_{1\leq i_{1}<\dots<i_{k}\leq n\>{}}\sum_{{}\>\{j_{1},\dots,j_{k}\}=\{i_{1},\dots,i_{k}\},j_{m}\neq i_{m}(1\leq m\leq k)}{\rm sgn}\left(\begin{array}[]{ccc}i_{1}&\cdots&i_{k}\\ j_{1}&\cdots&j_{k}\end{array}\right)\left(-q^{d_{i_{1}j_{1}}}\right)\dots\left(-q^{d_{i_{k}j_{k}}}\right).$ (2.3)
A permutation of $k$ letters such that no letters remain unchanged can be expressed as the product of several cyclic permutations of disjoint orbits with $k_{i}$ letters, where $\sum_{i}k_{i}=k$, and the expression is unique up to the order of the cycles. A cyclic permutation of $m$ letters $(i_{1},\dots,i_{m})$ corresponds to a single $m$-cycle $\Gamma_{i_{1}\dots i_{m}}=[x_{i_{1}},\dots,x_{i_{m}}]$. Since the signature of a cyclic permutation of $m$ letters is $(-1)^{m-1}$, we have
${\rm sgn}(i_{1},\dots,i_{m})\big{(}-q^{d_{i_{1}i_{2}}}\big{)}\dots\big{(}-q^{d_{i_{m-1}i_{m}}}\big{)}\big{(}-q^{d_{i_{m}i_{1}}}\big{)}=-q^{\ell(\Gamma_{i_{1}\dots i_{m}})}.$
Since a permutation of $k$ letters corresponds to a $k$-cycle in the sense of Definition 2.1, the right hand side of (2.3) is equal to
$\sum_{\Gamma\in\mathcal{C}_{k}(X)}{(-1)}^{c(\Gamma)}q^{\,\ell(\Gamma)},$
which, together with (2.3), implies the conclusion $\Box$
[…]
Corollary 2.3
From the generalized series expansion of the coefficients of the characteristic polynomial of the similarity matrix we obtain a multiset of edge lengths $[d_{ij}]_{i<j}$ and a multiset of lengths of $3$-cycles $[d_{ij}+d_{jk}+d_{ki}]_{i<j<k}$ by
$\sum_{i<j}q^{2d_{ij}}=-\tau_{2}(q),\quad\sum_{i<j<k}q^{d_{ij}+d_{jk}+d_{ki}}=\frac{1}{2}\tau_{3}(q).$
It follows that a complete graph and a three-point set can be identified by the characteristic polynomial.
Put $\mathcal{S}_{1}=[d_{ij}]_{i<j}$ and $\mathcal{S}_{3}=[d_{ij}+d_{jk}+d_{ki}]_{i<j<k}$.
Remark 2.4
(1) Let $P_{X}(t;\lambda)=\left|\lambda I-Z_{tX}\right|$ be the characteristic polynomial of the similarity matrix $Z_{tX}$ and $\tilde{\lambda}_{1}(t),\dots,\tilde{\lambda}_{n}(t)$ be the eigenvalues of $Z_{tX}$. Let $\bar{P}_{X}(t;\mu)=P_{X}(t;\mu+1)$ be a polynomial obtained from $P_{X}(t;\lambda)$ by setting $\mu=\lambda-1$, put $\tilde{\mu}_{i}(t)=\tilde{\lambda}_{i}(t)-1$ and let $\tilde{\tau}_{k}(t)$ $(1\leq k\leq n)$ be the $k$-th elementary symmetric polynomial of $\tilde{\mu}_{1}(t),\dots,\tilde{\mu}_{n}(t)$. Then
$\sum_{i<j}e^{-2d_{ij}t}=-\tilde{\tau}_{2}(t),\quad\sum_{i<j<k}e^{-(d_{ij}+d_{jk}+d_{ki})t}=\frac{1}{2}\tilde{\tau}_{3}(t),$
which implies
$\sum_{i<j}\,{(d_{ij})}^{m}=\frac{-1}{(-2)^{m}}\,\tilde{\tau}_{2}^{(m)}(0),\quad\sum_{i<j<k}\,{(d_{ij}+d_{jk}+d_{ki})}^{m}=\frac{(-1)^{m}}{2}\,\tilde{\tau}_{3}^{(m)}(0).$
The left hand sides above are the Newton polynomials of $d_{ij}$ and $d_{ij}+d_{jk}+d_{ki}$ respectively. Since the elementary symmetric polynomials can be given by the Newton polynomials, we obtain $[d_{ij}]_{i<j}$ from $\tilde{\tau}_{2}^{(m)}(0)$ $(1\leq m\leq{n\choose 2})$ and $[d_{ij}+d_{jk}+d_{ki}]_{i<j<k}$ from $\tilde{\tau}_{3}^{(m)}(0)$ $(1\leq m\leq{n\choose 3})$.
(2) Let us compare the eigenvalue information with the magnitude information.
The determinant and the sum of all the cofactors of $z_{X}(q)$ are both generalized polynomials, and since the determinant has the constant term $1$, it is invertible in the field of generalized rational functions. The sum of all the entries of $z_{X}(q)^{-1}$ is called the formal magnitude. Leinster showed in [5] Proposition 3.9 that the expansion of the formal magnitude in a “generalized formal power series” that allows non-integer exponents is given by
$m_{X}(q)=\sum_{k=0}^{\infty}{(-1)}^{k}\sum_{\begin{array}[]{c}\mbox{\scriptsize${i_{0}}\neq{i_{1}}\neq\dots\neq{i_{k}}$},\\[-2.56073pt] \mbox{\scriptsize$i_{0},i_{1},\dots,i_{k}\in\{1,\dots,n\}$}\end{array}}q^{d_{i_{0}i_{1}}+d_{i_{1}i_{2}}+\cdots+d_{i_{k-1}i_{k}}}.$
Thus in the case of formal magnitude, the coefficient of $q^{l}$ is given by the number of (not necessarily closed) paths with length $l$ together with the information of the number of points that gives the sign. On the other hand, (2.1) shows that the $k$-th elementary polynomial of the eigenvalues minus $1$ is given by the lengths of $k$-cycles together with the information of the number of connected components that gives the sign.
2.1 Generic case
A finite metric space $X=\{x_{1},\dots,x_{n}\}$ is generic if $d(x_{i},x_{j})$ $(i<j)$ are rationally independent ([8]). This condition was used to identify spaces by the formal magnitude, although with the characteristic polynomials a weaker assumption would be enough. Suppose the multisets of edge lengths $\mathcal{S}_{1}=[d_{ij}]_{i<j}$ and that of the lengths of $3$-cylcles $\mathcal{S}_{3}=[d_{ij}+d_{jk}+d_{ki}]_{i<j<k}$ are expressed as $\mathcal{S}_{1}=[a_{i}]_{i=1}^{N_{2}}$, where $N_{2}={n\choose 2}$, and $\mathcal{S}_{3}=[t_{j}]_{j=1}^{N_{3}}$, where $N_{3}={n\choose 3}$.
Definition 2.5
(cf. [8]) $X$ is weakly $3$-generic if $d_{ij}$ are mutually distinct and the sums of edge lengths do not match for different combinations of $3$ edges, namely,
[…]
has only the trivial solution.
In this case both multisets $\mathcal{S}_{1}$ and $\mathcal{S}_{3}$ become sets, which we denote by $S_{1}$ and $S_{3}$ in what follows in this Subsection.
We show that the class of weakly $3$-generic finite metric spaces is broader than the class of generic finite metric spaces. Roff and Yoshinaga gave the moduli space of unordered $n$-point metric spaces ([10] Section 2) as follows. Put
$\mathcal{L}_{n}=\left\{(\delta_{12},\delta_{13},\dots,\delta_{n-1\,n})\in(\mathbb{R}_{>0})^{N_{2}}\,\left|\,\begin{array}[]{l}\,\delta_{ij}+\delta_{jk}\geq\delta_{ik}\>\,\forall i,j,k\,(1\leq i,j,k\leq n),\\[1.42262pt] \mbox{where we put }\>\delta_{ii}=0,\,\delta_{ij}=\delta_{ji}\>\,\forall i,j\end{array}\right.\right\}.$
The symmetric group $\mathfrak{S}_{n}$ acts on $\mathcal{L}_{n}$ by $\sigma\cdot(\delta_{ij})=(\delta_{\sigma(i)\sigma(j)})$. Then the moduli space is given by $\mathcal{M}_{n}=\mathcal{L}_{n}/\mathfrak{S}_{n}$ equipped with the quotient topology.
The set of weakly $3$-generic finite metric spaces is open dense in $\mathcal{M}_{n}$ since the complement is a finite union of codimension one spaces. On the other hand, the set of generic finite metric spaces is not open dense in $\mathcal{M}_{n}$ since the complement includes $(\mathcal{L}_{n}\cap\mathbb{Q}^{N_{2}})/\mathfrak{S}_{n}$.
Theorem 2.6
A weakly $3$-generic finite metric space can be constructed from the characteristic polynomial of the similarity matrix of $tX=(X,td)$ $(t>0)$, or equivalently, from its eigenvalues.
In other words, both the characteristic polynomial and the eigenvalues are the complete invariants of weakly $3$-generic finite metric spaces.
Proof. We show inductively that a weakly $3$-generic finite metric space can be constructed from the sets ${S}_{1}$ and ${S}_{3}$. Remak that for each element $t_{m}\in S_{3}$ there is a unique triple $\{a_{i},a_{j},a_{k}\}\subset S_{1}$ that satisfies $t_{m}=a_{i}+a_{j}+a_{k}$.
(i) Take any element of $S_{3}$, say $t_{1}$. Let $\{a,b,c\}\subset S_{1}$ be the unique triple that satisfies $t_{1}=a+b+c$. Name the vertices $A,B$ and $C$ in the standard way, and choose the edge $BC$ with length $a$ as the base. Put $S_{3}(a)=\{t\in S_{3}\,|\,t=a+x+y,\,x,y\in S_{1}\}$. We may assume, by renumbering the subscripts of $t_{j}$ if necessary, that $S_{3}(a)=\{t_{1},\dots,t_{n-2}\}$.
(ii) Take $t_{2}$. Suppose $t_{2}=a+p+q$, where $p,q\in S_{1}\setminus\{a,b,c\}$. There are two ways to attach a triangle with edge lengths $a,p,q$, which we denote by $\triangle(a,p,q)$, to the edge $BC$. We choose one of them by attaching the edge with length $p$ to the vertex $B$ if there is $t\in S_{3}\setminus\{t_{1},t_{2}\}$ such that $t=p+c+x$ for some $x\in S_{1}\setminus\{a,b,c,p,q\}$, and to $C$ otherwise (Figure 1). Put $A_{2}$ to be the vertex of $\triangle(a,p,q)$ which is different from $B$ and $C$, then $AA_{2}=x$.
Figure 1:
(iii) Take $t_{3}\in S_{3}(a)$. Assume $t_{3}=a+r+s$, where $r,s\in S_{1}\setminus\{a,b,c,p,q\}$. We can determine the way how $\triangle(a,r,s)$ is attached to the edge $BC$ in the same way as above. Let $A_{3}$ be the vertex of $\triangle(a,r,s)$ which is different from $B$ and $C$. Then the length $AA_{3}$ is given by the unique element $u\in S_{1}\setminus\{a,b,c,p,q,r,s\}$ that satisfies $c+r+u\in S_{3}$ or $b+r+u\in S_{3}$ depending on the way how $\triangle(a,r,s)$ is attached to $BC$. The length $A_{2}A_{3}$ can be obtained in the same way.
(iv) Repeat the above procedure. $\Box$
We remark that the theorem does not exclude the possibility that a weakly $3$-generic space and a non weakly $3$-generic space have the same characteristic polynomial (or eigenvalues).
Without the assumption of genericity, the number of isometry classes of spaces with a given characteristic polynomial is bounded by a constant that depends only on the number of points, e.g. the factorial of (${{n}\choose 2}-1$.
2.2 Four-point sets
A four-point metric space $X=\{x_{1},x_{2},x_{3},x_{4}\}$ is identified by the formal magnitude if $X$ is generic or if $\max d_{ij}<2\min_{k\neq m}d_{kl}$ ([8] Theorem 2.4). We show that using the characteristic polynomials is stronger.
Theorem 2.7
A four-point metric space can be identified by the characteristic polynomial of the similarity matrix of $tX=(X,td)$ $(t>0)$, or equivalently, by its eigenvalues.
Proof. We use the $k$-th elementary symmetric polynomials $\tau_{k}(q)$ of $\mu_{k}(q)$ $(k=2,3,4)$, where $\mu_{i}(q)=\lambda_{i}(q)-1$. By Corollary 2.3, from $\tau_{2}(q)$ and $\tau_{3}(q)$ we obtain a multiset of edge lengths and a multiset of perimeters of four triangles, which we denote by $\mathcal{S}_{1}=[a,b,c,d,f,g]$ and $\mathcal{S}_{3}=[t_{1},t_{2},t_{3},t_{4}]$ respectively. Put $s$ to be the sum of six edge lengths $s=a+b+c+d+f+g$. The proof consists of two steps.
Step 1. First we show that we can obtain a multiset of the sums of the lengths of opposite (i.e. non-adjacent) edges, which we denote by $\mathcal{S}_{\rm opp}=[\alpha,\beta,\gamma]$. Note that $\alpha+\beta+\gamma=s$. We label the edge lengths $a,\dots,g$ as shown in Figure 2, and
Figure 2:
put
$\alpha=a+d,\>\>\beta=b+f,\>\>\gamma=c+g.$ (2.4)
Then the lengths of single $4$-cycles are given by
$\alpha+\beta=a+b+d+f,\>\>\beta+\gamma=b+c+f+g,\>\>\alpha+\gamma=a+c+d+f,$
[…]
$-2q^{\alpha+\beta}-2q^{\beta+\gamma}-2q^{\alpha+\gamma}+q^{2\alpha}+q^{2\beta}+q^{2\gamma}.$ (2.5)
There are following four cases (i) - (iv) according to the types of $\tau_{4}(q)$.
(1-i) $\tau_{4}(q)$ is of the form $-2\sum_{i=1}^{3}q^{A_{i}}+\sum_{i=j}^{3}q^{B_{j}},$ where $A_{i},B_{j}$ are mutually distinct. This case can occur if and only if $\alpha,\beta$ and $\gamma$ are mutually distinct and they do not form an arithmetic progression. In this case $\mathcal{S}_{\rm opp}$ is given by $\mathcal{S}_{\rm opp}=[B_{1}/2,B_{2}/2,B_{3}/2]$.
(1-ii) $\tau_{4}(q)$ is of the form
$-2q^{A_{1}^{\prime}}-2q^{A_{2}^{\prime}}-q^{A_{3}^{\prime}}+q^{B_{1}^{\prime}}+q^{B_{2}^{\prime}},$
where $A_{1}^{\prime},A_{2}^{\prime},A_{3}^{\prime},B_{1}^{\prime}$ and $B_{2}^{\prime}$ are mutually distinct. This case can occur if and only if $\alpha,\beta$ and $\gamma$ are mutually distinct and they form an arithmetic progression. We may assume without loss of generality that $\alpha+\beta=2\gamma$. Then cancellation between $-q^{\alpha+\beta}$ and $q^{2\gamma}$ occurs in (2.5), and $\mathcal{S}_{\rm opp}$ is given by $\mathcal{S}_{\rm opp}=[B_{1}^{\prime}/2,B_{2}^{\prime}/2,s-(B_{1}^{\prime}+B_{2}^{\prime})/2]$.
(1-iii) $\tau_{4}(q)$ is of the form $-4q^{A}+q^{B},$ where $A$ and $B$ are distinct. This case can occur if and only if exactly two of $\alpha,\beta$ and $\gamma$ coincide. We may assume without loss of generality that $\alpha=\beta\neq\gamma$. Then cancellation between $-2q^{\alpha+\beta}$ and $q^{2\alpha}+q^{2\beta}$ occurs in (2.5), and $\mathcal{S}_{\rm opp}$ is given by $\mathcal{S}_{\rm opp}=[B/2,(s-B/2)/2,(s-B/2)/2]$.
(1-iv) $\tau_{4}(q)$ is of the form $-3q^{A^{\prime}}$. This case can occur if and only if $\alpha,\beta$ and $\gamma$ are all equal. In this case $\mathcal{S}_{\rm opp}$ is given by $\mathcal{S}_{\rm opp}=[A^{\prime}/2,A^{\prime}/2,A^{\prime}/2]$.
Step 2. We show that if $X$ and $X^{\prime}$ both realize the same $\mathcal{S}_{1},\mathcal{S}_{3}$ and $\mathcal{S}_{\rm opp}$ then $X$ and $X^{\prime}$ are isometric. This is done by listing up all the possible configurations (i.e. the ways to assign elements of $\mathcal{S}_{1}$ to edges of a tetrahedron) with given data of $\mathcal{S}_{1}$ and $\mathcal{S}_{\rm opp}$ and then by showing that any two of them are isometric if they have the same $\mathcal{S}_{3}$.
Definition 2.8
By data of combination of opposite sides (abbreviated as dcos.) for $\mathcal{S}_{1}=[a,b,c,d,f,g]$ and $\mathcal{S}_{\rm opp}=[\alpha,\beta,\gamma]$ we mean a multiset $[\,[a,d],[b,f],[c,g]\,]$ that satisfies $[\alpha,\beta,\gamma]=[a+d,b+f,c+g]$. We denote it by $\mathcal{C}$.
For any dcos $\mathcal{C}$, there are at most two isometry classes that realize $\mathcal{C}$.
(2-i) Suppose both $X$ and $X^{\prime}$ with the same $\mathcal{S}_{1}$ and $\mathcal{S}_{\rm opp}$ have the same dcos. We may assume without loss of generality that it is given by $[\,[a,d],[b,f],[c,g]\,]$. Then there are two possibilities as illustrated in Figure 3 to be fixed to identify $X$.
Figure 3:
Let $\varphi_{i}=\varphi_{i}(a,b,c,d,f,g)$ (or $\psi_{j}$) $(1\leq i,j\leq 4)$ be the perimeters of four triangles of $X_{1}$ (or resp. $X_{2}$);
$\begin{array}[]{ll}\varphi_{1}(a,b,c,d,f,g)=a+b+c,&\quad\psi_{1}(a,b,c,d,f,g)=a+b+g,\\ \varphi_{2}(a,b,c,d,f,g)=a+f+g,&\quad\psi_{2}(a,b,c,d,f,g)=a+c+f,\\ \varphi_{3}(a,b,c,d,f,g)=b+d+g,&\quad\psi_{3}(a,b,c,d,f,g)=b+c+d,\\ \varphi_{4}(a,b,c,d,f,g)=c+d+f,&\quad\psi_{4}(a,b,c,d,f,g)=d+f+g.\end{array}$
Suppose $\mathcal{S}_{3}(X_{1})=\mathcal{S}_{3}(X_{2})$. There are $24$ possibilities for the system of linear equations obtained from $\mathcal{S}_{3}(X_{1})=\mathcal{S}_{3}(X_{2})$, each of which implies that $X_{1}$ is isometric to $X_{2}$. To be precise, for each permutation $\sigma\in\mathfrak{S}_{4}$ of four letters $\{1,2,3,4\}$, a system of linear equations
$\left\{\begin{array}[]{l}\varphi_{1}(a,b,c,d,f,g)=\psi_{\sigma(1)}(a,b,c,d,f,g),\\ \varphi_{2}(a,b,c,d,f,g)=\psi_{\sigma(2)}(a,b,c,d,f,g),\\ \varphi_{3}(a,b,c,d,f,g)=\psi_{\sigma(3)}(a,b,c,d,f,g),\\ \varphi_{4}(a,b,c,d,f,g)=\psi_{\sigma(4)}(a,b,c,d,f,g)\end{array}\right.$
implies that at least one of $a=d$, $b=f$ or $c=g$ holds [^3] The author used Maple 2023 for the computation of systems of linear equations in this subsection..
(2-ii) Suppose $X$ and $X^{\prime}$ with the same $\mathcal{S}_{1}$ and $\mathcal{S}_{\rm opp}$ have different dcos. Then, at least one of the elements of $\mathcal{S}_{\rm opp}$, $\alpha,\beta,\gamma$, say $\alpha$, is decomposed in two ways, say
$\alpha=a+b=(a-\varepsilon)+(b+\varepsilon)\qquad(\varepsilon\neq 0,\>a-\varepsilon\neq b).$
Let $c$ and $d$ be the remaining elements of $\mathcal{S}_{1}$, i.e., $\mathcal{S}_{1}=[a,b,a-\varepsilon,b+\varepsilon,c,d].$ Put $\delta=c+d$. Then $\delta\neq\alpha$ since if $\delta=\alpha$ then there is only one dcos, which is a contradiction.
Lemma 2.9
None of $X$ and $X^{\prime}$ has dcos of the form $\mathcal{C}_{\alpha\alpha\delta}=[[a,b],[a-\varepsilon,b+\varepsilon],[c,d]].$
Proof. Let $\mathcal{C}^{\prime}$ be another dcos;
$\mathcal{C}^{\prime}\colon\alpha=A_{11}+A_{12},\,\alpha=A_{21}+A_{22},\,\delta=B_{1}+B_{2}.$
Then at least two of $A_{ij}$ $(1\leq i,j\leq 2)$ belong to $\{a,b,a-\varepsilon,b+\varepsilon\}$. If the two are $A_{i1}$ and $A_{i2}$, as $[A_{3-i1},A_{3-i2}]$ cannot be $[b,c]$ because $\delta\neq\alpha$, $\mathcal{C}^{\prime}$ coincides with $\mathcal{C}_{\alpha\alpha\delta}$, which is a contradiction. If the two are $A_{1j}$ and $A_{2j^{\prime}}$ then $\mathcal{C}^{\prime}$ coincides with $\mathcal{C}_{\alpha\alpha\delta}$, which is a contradiction. $\Box$
Now we may assume without loss of generality $X$ has dcos
$\mathcal{C}_{\rm ii,0}\colon\alpha=(a-\varepsilon)+(b+\varepsilon),\,\beta=a+c,\,\gamma=b+d.$
Lemma 2.9, $c+d\neq\alpha$, $\varepsilon\neq 0$, and $a-\varepsilon\neq b$ imply that the dcos of $X^{\prime}$ is either
$\mathcal{C}_{\rm ii,1}\colon\alpha=a+b,\,\beta=(a-\varepsilon)+d,\,\gamma=(b+\varepsilon)+c$
or
$\mathcal{C}_{\rm ii,2}\colon\alpha=a+b,\,\beta=(b+\varepsilon)+d,\,\gamma=(a-\varepsilon)+c.$
(2-ii-1) Suppose $X^{\prime}$ has dcos $\mathcal{C}_{\rm ii,1}$. Then $\beta=a+c=a+d-\varepsilon$, which implies that $d=c+\varepsilon$, and hence $\mathcal{C}_{\rm ii,0}$ and $\mathcal{C}_{\rm ii,1}$ are given by
$\begin{array}[]{lll}\mathcal{C}_{\rm ii,0}\colon\alpha=(a-\varepsilon)+(b+\varepsilon),&\,\beta=a+c,&\,\gamma=b+(c+\varepsilon),\\ \mathcal{C}_{\rm ii,1}\colon\alpha=a+b,&\,\beta=(a-\varepsilon)+(c+\varepsilon),&\,\gamma=(b+\varepsilon)+c.\end{array}$
There are two realizations of $\mathcal{C}_{\rm ii,0}$, $X_{01}$ and $X_{02}$ (Figure 5), and two realizations of $\mathcal{C}_{\rm ii,1}$, $X_{11}$ and $X_{12}$ (Figure 5).
Figure 4:
Figure 5:
Suppose $X_{0i}$ and $X_{1j}$ $(i,j=1,2)$ have the same $\mathcal{C}_{3}$. Then at least one of the systems of four linear equations must be satisfied. Examine all the $4\times 4!=96$ systems of linear equations, and we obtain either the isometry between $X_{0i}$ and $X_{1j}$ or $\varepsilon=0$, which contradicts our assumption (2-ii). Let us give two examples. Assume $X_{01}$ and $X_{11}$ have the same $\mathcal{C}_{3}$. Put
$\begin{array}[]{ll}\varphi_{1}(a,b,c,\varepsilon)=2a+c,&\quad\psi_{1}(a,b,c,\varepsilon)=2a+b,\\ \varphi_{2}(a,b,c,\varepsilon)=a+b+c-\varepsilon,&\quad\psi_{2}(a,b,c,\varepsilon)=a+b+c-\varepsilon,\\ \varphi_{3}(a,b,c,\varepsilon)=a+2b+\varepsilon,&\quad\psi_{3}(a,b,c,\varepsilon)=a+2c+\varepsilon,\\ \varphi_{4}(a,b,c,\varepsilon)=b+2c+2\varepsilon,&\quad\psi_{4}(a,b,c,\varepsilon)=2b+c+2\varepsilon.\end{array}$
$\bullet$ The system of linear equations $\varphi_{1}=\psi_{3},\varphi_{2}=\psi_{2},\varphi_{3}=\psi_{4},\varphi_{4}=\psi_{1}$, i.e.
$\left\{\begin{array}[]{rl}2a+c&=a+2c+\varepsilon,\\ a+b+c-\varepsilon&=a+b+c-\varepsilon,\\ a+2b+\varepsilon&=2b+c+2\varepsilon,\\ b+2c+2\varepsilon&=2a+b\end{array}\right.$
has a solution $a=c+\varepsilon$. Then $X_{01}$ and $X_{11}$ are as illustrated in Figure 6, and they are isometric.
Figure 6:
$\bullet$ The system of linear equations $\varphi_{1}=\psi_{1},\varphi_{2}=\psi_{3},\varphi_{3}=\psi_{2},\varphi_{4}=\psi_{4}$, i.e.
$\left\{\begin{array}[]{rl}2a+c&=2a+b,\\ a+b+c-\varepsilon&=a+2c+\varepsilon,\\ a+2b+\varepsilon&=a+b+c-\varepsilon,\\ b+2c+2\varepsilon&=2b+c+2\varepsilon\end{array}\right.$
has a solution $b=c$ and $\varepsilon=0$, which contradicts our assumption that $\varepsilon\neq 0$. $\Box$
Since there is a pair of non-isometric four-point spaces with the same magnitude (homology groups) as illustrated in Figure 7, we have
Corollary 2.10
The characteristic polynomial, or equivalently, the eigenvalues of the similarity matrix of $tX$ are not weaker than the magnitude homology groups and hence the magnitude in classifying finite metric spaces.
3 Non-isometric spaces with the same magnitude
Let us examine some examples to what extent the eigenvalues identify non-generic finite metric spaces. Several methods are known to construct a pair of non-isometric spaces with the same magnitude or even the isomorphic magnitude homology groups (1-4 below). We show that they have different characteristic polynomials. The author used Maple 2023 for the computation in this section. The difference in the characteristic polynomials can be explained by the difference in the cycle status (cd. Lemma 2.2 and Corollary 2.3).
(1) Suppose $X=X_{1}\cup X_{2}$, $X_{1}\cap X_{2}=\{z\}$ and $d(x,y)=d(x,z)+d(z,y)$ for any $x\in X_{1}$ and $y\in X_{2}$. Then $X$ is called the wedge sum of $X_{1}$ and $X_{2}$, denoted by $X_{1}\vee_{z}X_{2}$. The magnitude of $X$ can be obtained from those of $X_{1}$ and $X_{2}$ ([4] Corollary 2.3.3), and the same statement holds for the magnitude homology groups, which can be verified by applying the Mayer-Vietoris exact sequence.
[…]
Two graphs in Figure 7 (Leinster [4] Example 2.3.5), both of which are the wedge sums of line graphs with two and three vertices, have the same magnitude and magnitude homology groups, but different characteristic polynomials.
$\begin{array}[]{rcl}p_{{\rm left}}(q;\lambda)&=&\displaystyle\lambda^{4}-4\lambda^{3}-\left(q^{6}+2q^{4}+3q^{2}-6\right)\lambda^{2}-2\left(q^{6}-3q^{2}+2\right)\lambda-q^{6}+3q^{4}-3q^{2}+1,\\[2.84526pt] p_{{\rm right}}(q;\lambda)&=&\displaystyle\lambda^{4}-4\lambda^{3}-\left(3q^{4}+3q^{2}-6\right)\lambda^{2}-2\left(q^{6}-3q^{2}+2\right)\lambda-q^{6}+3q^{4}-3q^{2}+1.\end{array}$
The difference of the coefficients of $\lambda^{2}$ comes from the difference of the number of pairs of points with distance $2$ and $3$.
By taking the wedge sum of three spaces in different ways, we can easily create examples of spaces where the magnitude homology group is isomorphic but not isometric (cf. Figure 8).
Figure 8: Two wedge sums of three cycle graphs
Two graphs in Figure 8 have different characteristic polynomials since they have different diameters (cf. Corollary 2.3).
(2) Let $E$ and $B$ be metric spaces and $\pi\colon E\to B$ be a distance-decreasing map. Suppose for any $a\in E$ and $b^{\prime}\in B$ there exists $a_{b^{\prime}}\in\pi^{-1}(b^{\prime})$ such that for any $a^{\prime}\in\pi^{-1}(b^{\prime})$ there holds $d(a,a^{\prime})=d(\pi(a),b^{\prime})+d(a_{b^{\prime}},a^{\prime})$. Then $\pi^{-1}(b)$ and $\pi^{-1}(b^{\prime})$ are isometric for any $b,b^{\prime}\in B$ ([4] Lemma 2.3.10). $\pi\colon E\to B$ is called the metric fibration with fiber $F=\pi^{-1}(b)$ ([4] Definition 2.3.8). The magnitude of $E$ can be obtained from those of $B$ and $F$ ([4] Theorem 2.3.11), and the same statement holds for the magnitude homology groups ([1] Corollary 2.17).
Figure 9: Metric fibration
The characteristic polynomials of the two graphs in Figure 9 are given by
$\begin{array}[]{rcl}p_{{\rm left}}(q;\lambda)&=&\displaystyle\lambda^{6}-6\lambda^{5}-3\left(2\,q^{4}+3\,q^{2}-5\right)\lambda^{4}-4\left(3\,q^{5}+q^{3}-9q^{2}+5\right)\lambda^{3}+\dots,\\[5.69054pt] p_{{\rm right}}(q;\lambda)&=&\displaystyle\lambda^{6}-6\lambda^{5}-3\left(2\,q^{4}+3\,q^{2}-5\right)\lambda^{4}-4\left(q^{6}+3\,q^{4}-9q^{2}+5\right)\lambda^{3}+\dots.\end{array}$
Note that the maximal length of the $3$-cycles is $5$ in the left graph, whereas it is $6$ in the right graph, which causes the difference of the coefficients of $\lambda^{5}$.
(3) Let $G$ and $H$ be graphs, $g_{+},g_{-}$ and $h_{+},h_{-}$ be two distinct vertices of $G$ and $H$ respectively. Let $X$ (or $Y$) be a graph obtained from the disjoint union of $G$ and $H$ by identifying $g_{+}$ with $h_{+}$ and $g_{-}$ with $h_{-}$ (or respectively, $g_{+}$ with $h_{-}$ and $g_{-}$ with $h_{+}$), and removing a double edge if it is created in the gluing procedure. The graphs $X$ and $Y$ are said to differ by a Whitney twist. Leinster showed that if the gluing points are adjacent, i.e. if there is an edge joining them, then $|X|=|Y|$ ([5] Theorem 5.2).
Figure 10: Two graphs that differ by a Whitney twist
Let us consider two graphs in Figure 10. Although the gluing points are not adjacent and hence we cannot apply Leinster’s result, these two graphs have isomorphic magnitude homology groups. This is a consequence of Mayer-Vietoris for magnitude homology ([3] Theorem 6.6). In fact, let $X$ be either of the graphs in Figure 10, if we set $G$ to be the union of the triangle and the square and $H$ the remaining square, then $(X;G,H)$ is a projecting decomposition of [3] Definition 6.3.
The characteristic polynomials of two graphs in Figure 10 are given by
$\begin{array}[]{rcl}p_{{\rm left}}(q;\lambda)&=&\displaystyle\lambda^{7}-7\lambda^{6}-\left(3\,q^{6}+9\,q^{4}+9\,q^{2}-21\right)\lambda^{5}+\dots,\\[5.69054pt] p_{{\rm right}}(q;\lambda)&=&\displaystyle\lambda^{7}-7\lambda^{6}-\left(4\,q^{6}+8\,q^{4}+9\,q^{2}-21\right)\lambda^{5}+\dots,\end{array}$
The difference of the coefficients of $\lambda^{5}$ comes from the difference of the number of pairs of points with distance $2$ and $3$.
Next, two graphs in Figure 11 have the same magnitude by Leinster’s theorem.
Figure 11: Another example of Whitney twist
The characteristic polynomials of two graphs in Figure 11 are given by
$\begin{array}[]{rcl}p_{{\rm left}}(q;\lambda)&=&\displaystyle\lambda^{10}-10\lambda^{9}-\left(q^{10}+5\,q^{8}+10\,q^{6}+16\,q^{4}+13\,q^{2}-45\right)\lambda^{8}+\dots,\\[5.69054pt] p_{{\rm right}}(q;\lambda)&=&\displaystyle\lambda^{10}-10\lambda^{9}-\left(6\,q^{8}+10\,q^{6}+16\,q^{4}+13\,q^{2}-45\right)\lambda^{8}+\dots.\end{array}$
The difference of the coefficients of $\lambda^{8}$ comes from the difference of the number of pairs of points with distance $4$ and $5$.
(4) The notion of isomer was introduced in [9] to construct a space that has the same magnitude with, but is not isometric with the regular $n$-gon or an $n$-cycle graph. To be precise, an isomer of a regular $n$-gon can be constructed for $n=6$ and $n\geq 8$, whereas an isomer of an $n$-cycle graph can only be constructed for an even number greater than or equal to $6$. The magnitude homology groups of the original space and the isomer given in [9] are isomorphic in the case of regular polygons, but not isomorphic in the case of cycle graphs.
Let us consider the case when $n=6$. The construction of the isomer is shown in Figure 12.
Figure 12: Thin single lines represent length $d_{1}$, double lines represent length $d_{2}$, and dotted lines represent length $d_{3}$. They represent a regular hexagon $R_{6}$ and its isomer $R_{6}^{\prime}$ when $(d_{1},d_{2},d_{3})=(1,\sqrt{3},2)$, and a $6$-cycle graph $C_{6}$ and its isomer $C_{6}^{\prime}$ when $(d_{1},d_{2},d_{3})=(1,2,3)$.
Regular hexagon and its isomer have different characteristic polynomials since they have different maximal lengths of $3$-cycles. In fact we have
$\begin{array}[]{rcl}p_{R_{6}}(q;\lambda)&=&\lambda^{6}-6\lambda^{5}-3\left(q^{4}+2\,q^{2\sqrt{3}}+2\,q^{2}+5\right)\lambda^{4}\\[5.69054pt] &&+2\left(-2\,q^{3\sqrt{3}}-12\,q^{(3+\sqrt{3}\,)}-6\,q^{(2+\sqrt{3}\,)}+6\,q^{4}+12\,q^{2\sqrt{3}}+12\,q^{2}-10\right)\lambda^{3}+\cdots,\\[8.53581pt] p_{R_{6}^{\prime}}(q;\lambda)&=&\lambda^{6}-6\lambda^{5}-3\left(q^{4}+2\,q^{2\sqrt{3}}+2\,q^{2}+5\right)\lambda^{4}\\[5.69054pt] &&+2\left(-12\,q^{(3+\sqrt{3}\,)}-6\,q^{(1+2\sqrt{3}\,)}+6\,q^{4}+12\,q^{2\sqrt{3}}-2\,q^{3}+12\,q^{2}-10\right)\lambda^{3}+\cdots.\\[8.53581pt] \end{array}$
$6$-cycle graph and its isomer have different characteristic polynomials since the number of $3$-cycles with maximal length $6$ is different. In fact we have
$\begin{array}[]{rcl}p_{C_{6}}(q;\lambda)&=&\lambda^{6}-6\lambda^{5}-3\left(q^{3}+2\,q^{4}+2\,q^{2}-5\right)\lambda^{4}+4\left(-4\,q^{6}+3\,q^{4}+6\,q^{2}-5\right)\lambda^{3}+\dots,\\[8.53581pt] p_{C_{6}^{\prime}}(q;\lambda)&=&\lambda^{6}-6\lambda^{5}-3\left(q^{3}+2\,q^{4}+2\,q^{2}-5\right)\lambda^{4}+4\left(-3\,q^{6}-3\,q^{5}+6\,q^{4}-q^{3}+6\,q^{2}-5\right)\lambda^{3}+\dots.\end{array}$
When $n=8$, the construction of isomers is illustrated in Figure 13.
Figure 13: Thin single lines, double lines, dotted lines and thick lines represent lengths $d_{1},d_{2},d_{3}$ and $d_{4}$ respectively. The figure represents an isomer of a regular octagon if $d_{i}=\sin(\pi i/8)/\sin(\pi/8)$ and an isomer of a $8$-cycle graph if $d_{i}=i$.
The regular octagon $R_{8}$ and its isomer $R_{8}^{\prime}$ have different characteristic polynomials. In fact they have the same coefficients of $\lambda^{8},\lambda^{7},\lambda^{6}$ and $\lambda^{5}$, but different coefficients of $\lambda^{4}$ because the numbers of $4$-cycles with length $4$ are different. The same statement holds for the $8$-cycle graph and its isomer.
When $n=9$, by putting $|m|_{9}=\min\{|m|,9-|m|\}$ for $1\leq m\leq 9$, $\delta_{m}=(\sin(m\pi/9))/(\sin(\pi/9))$ and $d^{\prime}_{ij}=\delta_{|j-i|_{9}}$ if $|j-i|_{9}\leq 2$, $d^{\prime}_{ij}=\delta_{4}$ if $|j-i|_{9}=3$ and $d^{\prime}_{ij}=\delta_{3}$ if $|j-i|_{9}=4$, we obtain an isomer $R_{9}^{\prime}$ of a regular nonagon $R_{9}$. Then $R_{9}$ and $R_{9}^{\prime}$ have different characteristic polynomials since the maximal lengths of $3$-cycles are different; $R_{9}$ has $3\delta_{3}$ whereas $R_{9}^{\prime}$ has $3\delta_{4}$.
(5) Let $\Sigma=\{1,2,\dots,16\}$ be a metric space of $16$ points equally placed on a circle with a circumference of $16$ in the cyclic order, and the distance between them is determined by the arc-length on $\Sigma$. Let $\Sigma_{i}$ $(i=1,\dots,6)$ be metric subspaces of $\Sigma$ as illustrated in Figures 14 and 15. The convex hulls of $\Sigma_{1}$ and $\Sigma_{2}$ are known as Mallows and Clark’s counter example to Blaschke’s problem whether a space is characterized by the chord length distribution ([7]).
Figure 14:
Figure 15:
Direct computation shows
$m_{\Sigma_{1}}(q)=m_{\Sigma_{2}}(q)=m_{\Sigma_{3}}(q)\neq m_{\Sigma_{4}}(q)\neq m_{\Sigma_{5}}(q)=m_{\Sigma_{6}}(q)\neq m_{\Sigma_{1}}(q),$
and the characteristic polynomials of $Z_{t\Sigma_{i}}$ are all mutually distinct. In the case of $\Sigma_{1}$ and $\Sigma_{2}$ the difference in the characteristic polynomials appears first in the coefficients of $\lambda^{9}$, otherwise in the coefficients of $\lambda^{10}$.
4 Cospectral graphs
Two non-isomorphic graphs are said to be cospectral or isospectral if the adjacency matrices have the same multisets of eigenvalues. Figures 17 and 17 show two examples of pairs of cospectral graphs that have different characteristic polynomials of similarity matrices and different magnitudes.
Figure 16:
Figure 17:
On the other hand, two cospectral graphs in Figure 18 have the same characteristic polynomials of the similarity matrices and the same magnitudes.
Figure 18: Isospectral enneahedra (picture taken from wiki)